# Linear Attention Based Deep Nonlocal Means Filtering for Multiplicative Noise Removal

### v1

@@ Abstract
Multiplicative noise widely exists in radar images, medical images and other important fields’ images. Compared to normal noises, multiplicative noise has a generally stronger effect on the visual expression of images. Aiming at the denoising problem of multiplicative noise, we linearize the nonlocal means algorithm with deep learning and propose a linear attention mechanism based deep nonlocal means filtering (LDNLM). Starting from the traditional nonlocal means filtering, we employ deep channel convolution neural networks to extract the information of the neighborhood matrix and obtain representation vectors of every pixel. Then we replace the similarity calculation and weighted averaging processes with the inner operations of the attention mechanism. To reduce the computational overhead, through the formula of similarity calculation and weighted averaging, we derive a nonlocal filter with linear complexity. Experiments on both simulated and real multiplicative noise demonstrate that the LDNLM is more competitive compared with the state-of-the-art methods. Additionally, we prove that the LDNLM possesses interpretability close to traditional NLM.
Keywords:
Image denoising Non-local means Multiplicative noise
1 Introduction
@@
Additionally, since the deep neural network is only a black box in the denoising model, for the above-mentioned deep learning-based image denoising methods, it is difficult to explain how the results of them are produced. To address the above problem, [23], [7] and [11] attempted to incorporates the deep models with NLM, but the interpretability of them is still limited. Unlike them, we employ deep channel CNN to extract the pixels’ semantics first. Then, the inner product of the extracted vectors is adopted to calculate the similarities of pixels. Finally, the weighted averaging of extracted vectors and a linear projection are performed to obtain the filtered results. Additionally, the kernel functions are used to make the similarity calculation linearized. The main contributions of this work can be summarized as follows:
@@
  • By changing the calculation order of inner vectors, we derive a nonlocal means denoising algorithm with linear complexity.
@@
2 Related Work
2.1 Nonlocal Denoising
@@ 3 Method
3.1 Deep Nonlocal Means Filtering
@@
3.2 Linear Attention Based Deep Nonlocal Means Filtering
@@
The fundamental reason for the quadratic complexity lies in the Eq. 7. Where all $N$ pixels in the search window should be used to calculate similarity with all other $N-1$ pixels. For the resulting vector $V_{i}^{\prime}$ of $i$-th pixel, we can write a generalized attention equation for any similarity function as follows:
@@ 4 Experiments
We compare the proposed method with traditional methods and state-of-the-art deep learning-based methods in the experiments. Specifically, the methods for comparison include non-learning based methods nonlocal means [3], BM3D[8], supervised denoising method SAR-CNN[38], MONet [35], SAR-CAM[20] and CNN-NLM [7].
@@ 4.1 On Simulation Images
The average PSNR and SSIM of each model are as Tab. 1, and the denoising results of 2 example simulated images are shown in Fig. 2.
Figure 2: Performance of each model on example simulated images containing multiplicative noise. From left to right: Noisy, Reference, NLM, BM3D, SAR-CNN, MONet, SAR-CAM and LDNLM. From top to bottom: dense residential and parking lot.
Table 1: Average Performance of Each Model on Simulated SAR Images.
@@ 4.2 On Real SAR Images
Figure 3: Performance of each model on real SAR images: TerraSAR-Mountain. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM.
Figure 4: Performance of each model on real SAR images: TerraSAR-Urban. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM.
Figure 5: Performance of each model on real SAR images: The patch indicated by the red boxes in Fig. 4. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM.
Figure 6: The ratio image of each model on SAR images: TerraSAR-Urban. From left (top) to right (bottom): NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM. The ratio image follows the statistical properties of the gamma noise ideally.
The selected mountain SAR image and the filtered results of each model are shown in Fig. 3, and the selected urban SAR image and the filtered results of each model are shown in Fig. 4. The ENL and $\mathcal{M}$ of all filtered results are shown as Tab. 2, the blue boxes in the figures indicate the homogeneous regions bounded for the calculation of ENL. To better compare the results of each model, we enlarge the part bounded by the red box in Fig. 4 as Fig. 5. Additionally, Fig. 6 shows the ratio images for each model.
Table 2: Performance of Each Model on Real SAR Images. Optimal values are bolded and the second best values are italicized.
@@
The ratio image of LDNLM is nearly pure noise, while others’ contain many structures of streets. We can observe that the LDNLM gets the relative best texture detail repairing performance in two common SAR scenes. Compared to other methods, LDNLM achieves a competitive noise removal degree. Combing the ratio images and enlarged images of urban scenes, we can find that LDNLM is able to repair structures like roads and buildings whereas SAR-CNN and MONet are doing not good. It can be concluded that the LDNLM is better at balancing multiplicative noise removal and structure restoration.
4.3 Ablation Study
@@
Tab. 3 shows the results on the simulated noisy image same with Sec. 4.1.
Table 3: Performance of Each Model on Real SAR Images.

### v2

@@ Abstract
Multiplicative noise widely exists in radar images, medical images and other important fields’ images. Compared to normal noises, multiplicative noise has a generally stronger effect on the visual expression of images. Aiming at the denoising problem of multiplicative noise, we linearize the nonlocal means algorithm with deep learning and propose a linear attention mechanism based deep nonlocal means filtering (LDNLM). Starting from the traditional nonlocal means filtering, we employ deep channel convolution neural networks to extract the information of the neighborhood matrix and obtain representation vectors of every pixel. Then we replace the similarity calculation and weighted averaging processes with the inner operations of the attention mechanism. To reduce the computational overhead, through the formula of similarity calculation and weighted averaging, we derive a nonlocal filter with linear complexity. Experiments on both simulated and real multiplicative noise demonstrate that the LDNLM is more competitive compared with the state-of-the-art methods. Additionally, we prove that the LDNLM possesses interpretability close to traditional NLM. The source code and pre-trained model are available at https://github.com/ShowiBin/LDNLM.
I Introduction
@@
Additionally, since the deep neural network is only a black box in the denoising model, for the above-mentioned deep learning-based image denoising methods, it is difficult to explain how the results of them are produced. To address the above problem, [23], [7], [11] and [41] attempted to incorporates the deep models with NLM, but the interpretability of them is still limited. Unlike them, we employ deep channel CNN to extract the pixels’ semantics first. Then, the inner product of the extracted vectors is adopted to calculate the similarities of pixels. Finally, the weighted averaging of extracted vectors and a linear projection are performed to obtain the filtered results. Additionally, the kernel functions are used to make the similarity calculation linearized. The main contributions of this work can be summarized as follows:
@@
  • Through changing the calculation order of inner vectors, we derive a nonlocal means denoising algorithm with linear complexity.
@@
II RELATED WORK
II-A Nonlocal Denoising
@@ III METHODOLOGY
III-A Deep Nonlocal Means Filtering
@@
III-B Linear Attention Based Deep Nonlocal Means Filtering
@@
The fundamental reason for the quadratic complexity lies in the eq. 7. Where all $N$ pixels in the search window should be used to calculate similarity with all other N1N-1 pixels. For the resulting vector $V_{i}^{\prime}$ of $i$-th pixel, we can write a generalized attention equation for any similarity function as follows:
@@ IV Experiments
We compare the proposed method with traditional methods and state-of-the-art deep learning-based methods in the experiments. Specifically, the methods for comparison include non-learning based methods nonlocal means [3], BM3D[8], supervised deep learning method SAR-CNN[38], MONet [35], SAR-CAM[20], CNN-NLM [7] and Trans-SAR[31], unnsupervised deep learning method in [40].
@@ IV-A On Simulation Images
The average PSNR and SSIM of each model are as table I. Additionally, we select NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM, get their denoising results of 2 example simulated images, as fig. 2.
Figure 2: Performance of each model on example simulated images containing multiplicative noise. From left to right: Noisy, Reference, NLM, BM3D, SAR-CNN, MONet, SAR-CAM, CNN-NLM and LDNLM. From top to bottom: dense residential and parking lot.
TABLE I: Average Performance of Each Model on Simulated SAR Images.
@@ IV-B On Real SAR Images
Figure 3: Performance of each model on real SAR images: TerraSAR-Mountain. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, method in [40], CNN-NLM and LDNLM.
Figure 4: Performance of each model on real SAR images: TerraSAR-Urban. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, method in [40], CNN-NLM and LDNLM.
Figure 5: Performance of each model on real SAR images: The patch indicated by the red boxes in fig. 4. From left (top) to right (bottom): Noisy, NLM, BM3D, SAR-CNN, MONet, method in [40], CNN-NLM and LDNLM.
Figure 6: The ratio image of each model on SAR images: TerraSAR-Urban. From left (top) to right (bottom): NLM, BM3D, SAR-CNN, MONet, method in [40], CNN-NLM and LDNLM. The ratio image follows the statistical properties of the gamma noise ideally.
To compare the denosing results intuitively, we select NLM, BM3D, SAR-CNN, MONet, method in [40], CNN-NLM and LDNLM, and then get their denoising results of sample SAR images. The selected mountain SAR image and the filtered results of each model are shown in fig. 3, and the selected urban SAR image and the filtered results of each model are shown in fig. 4. The ENL and $\mathcal{M}$ of all filtered results are shown as table II, the blue boxes in the figures indicate the homogeneous regions bounded for the calculation of ENL. To better compare the results of each model, we enlarge the part bounded by the red box in fig. 4 as fig. 5. Additionally, fig. 6 shows the ratio images for each model.
TABLE II: Performance of Each Model on Real SAR Images. Optimal values are bolded and the second best values are italicized.
@@
The ratio image of LDNLM is nearly pure noise, while others’ contain many structures of streets. We can observe that the LDNLM gets the relative best texture detail repairing performance in two common SAR scenes. Compared to other methods, LDNLM achieves a competitive noise removal degree. Combing the ratio images and enlarged images of urban scenes, we can find that LDNLM is able to repair structures like roads and buildings whereas SAR-CNN and MONet are doing not good. Unsupervised method in [40] achieves a remarkable performance on speckle removal with it’s great generating ability, but it does not perform well in texture preservation. It can be concluded that the LDNLM is better at balancing multiplicative noise removal and structure restoration.
IV-C Ablation Study
@@
table III shows the results on the simulated noisy image same with section IV-A.
TABLE III: Performance of Each Model on Real SAR Images.